OpenStudy (anonymous):

Find sin (x+y), assuming that sin x= -2/3, cos y=1/4 and both x and y are in quadrant IV.

OpenStudy (anonymous):

sin(x+y)=sinxcosy+cosysinx

OpenStudy (anonymous):

are you familiar with this identity?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have to look at my notes

OpenStudy (anonymous):

do you know how to find cos(x) if you are given sin(x)? because you need that too.

OpenStudy (anonymous):

sin x=1/cos x

OpenStudy (anonymous):

This is incorrect I think you are mistaking it with sec(x)=1/cos (x)

OpenStudy (anonymous):

oh, you're right

OpenStudy (anonymous):

Okay, so we know this identity sin(x+y)=sin(x)cos(y)+cos(y)sin(x) now we have to plug in what we know

OpenStudy (anonymous):

You know sin(x),cos(y) you don't know sin(y),cos(x)

OpenStudy (anonymous):

i have no idea what to do

OpenStudy (anonymous):

So you need to find what you don't know are you familiar with soh-cah-toa

OpenStudy (anonymous):

a bit, but I don't even know how you got sin (x+y)=sin (x)cos(y)+cos(y)sin(x)

OpenStudy (anonymous):

this is identity, you just need to know that

OpenStudy (anonymous):

ok, I see it in my notes under cosine of a sum or difference